OpenStudy (anonymous):

Nathan took his car in for service and repairs. He had a coupon for 15% off. Original Prices • Labor $288.85 • Parts $329.62 • Other $71.49 Which amount is closest to Nathan's total costs after the 15% discount and including 6% sales tax? [Assume tax is applied after the discount is applied.]

OpenStudy (cj49):

288.85+329.62+71.49=689.96 689.96=689.96x15/100=103.94 689.96-103.64=586.32 586.32x6/100=35.17 586.32+35.17=621.49